# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v4

@@ Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states. These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling. We demonstrate our approach in two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms. Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study. Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
@@ 1 Introduction
Understanding molecular conformational changes is fundamental to understanding biological processes, from enzyme catalysis to protein folding. Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes. This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935). As a result, direct simulation methods spend the vast majority of computational effort sampling stable states, making the observation of transitions between them prohibitively expensive.
@@
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths through a discrete series of molecular configurations, connected by artificial spring forces to ensure continuity or as splines. Although grounded in solid theoretical principles, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for many small-scale systems, it has not yet fully benefited from modern optimization approaches that could address these challenges.
@@
Concurrent advances in scientific machine learning and molecular simulation have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical forcefield models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives needed for path optimization.
@@
The last decades have seen numerous approaches emerge to tackle the transition path sampling problem. Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample full transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024). Notably, Du et al. (2024) use a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble via Doob’s h-transform. This approach provides rich statistical information about transition mechanisms. Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of minimum-energy paths inspired by Nudged Elastic Band (NEB), incorporating specialized components for tangential forces, path velocity, and transition state energetics. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a forcefield within a chain-of-states framework to generate plausible transition distributions. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to large biomolecular systems with explicit solvent while maintaining fast training times.
Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy path finding problem using a continuous optimization framework. Our approach leverages the universal function approximation capabilities of neural networks to represent conformational transitions as differentiable mappings from a progress coordinate to molecular configurations. By embedding physical principles directly into the optimization objective through a physics-informed neural network (PINN) framework, we eliminate the need for explicit reparameterization steps that burden traditional methods. This design enables applications to complex biomolecular systems in minutes on standard hardware while extracting physically meaningful transition mechanisms. The resulting pathways can serve as valuable starting points for established transition path sampling methods (Dellago et al., 1998; Bolhuis et al., 2002) when exploration of the complete ensemble is desired.
@@ 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To derive our loss function, which when minimized yields a continuous curve representing the minimum-energy path connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods. However, we further simplify the approach after reaching the geometric formulation. We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
The exponent term can be recognized as the Onsager-Machlup action functional Onsager & Machlup (1953):
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. To further simplify, we expand the integrand of the action functional:
@@
Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization. We can therefore focus on minimizing the remaining terms. Applying the Cauchy-Schwarz inequality:
This upper bound becomes an equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape. Crucially, in our formulation, we allow the path to optimize its shape and effective length. This freedom enables the path to adjust to follow the energy gradient (or its opposite), making the above upper bound tight at the optimum. When we abandon fixed-time parameterization and allow the path to stretch as needed to align with energy gradients, we can transform the time integral into a purely geometric one by defining the arc-length element along the spatial curve :
[…]
@@
For a path segment between adjacent points, applying Taylor’s theorem, we can express the energy difference as:
Where is the angle between the gradient and displacement vectors. For minimum-energy paths, the displacement aligns with or against the gradient direction () giving:
@@
Since the latter terms are constants during optimization, this leads to our final loss function. To implement this in practice, we sample B points along a parametrized path and directly minimize the sum of energies at these points, yielding:
where are uniformly or randomly sampled points in . A complete derivation with all mathematical details is provided in Appendix A.1. By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
2.1.1 Neural Representation of Minimum-Energy Paths
We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms:
Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
@@ 2.1.2 Path Optimization
We minimize directly through stochastic gradient descent. For each batch of points along the path:
Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary: the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
2.2 Established Approaches to Minimum-Energy Path Finding
Classical approaches to minimum-energy path discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation. This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization. The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as "Transition-Path Theory and Path-Finding Algorithms for the Study of Rare Events" by E & Vanden-Eijnden (2010). These established techniques illustrate both foundational concepts and limitations that our approach addresses.
@@ 2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized by that connects these states. The MEP satisfies:
@@
Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all . In practice, the string is discretized into images and evolved through a two-step procedure:
@@ 2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition . This method requires explicit reparameterization steps that can introduce computational inefficiencies.
@@ 2.2.2 Chain-of-States Methods
An alternative perspective views MEP finding as optimization of a chain of states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008). Given a sequence with and , the objective function is defined as:
where and is a spring constant. The spring term enforces the continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost. The neural representation avoids this limitation through its natural continuity properties.
@@ 3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method with the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
@@ 3.1 AIB9 Transition Path Discovery
We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods. This system effectively bridges abstract models and complex proteins: it is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes; its energy landscape is well-characterized; and its transitions are easily visualized through dihedral angle projections. Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields. We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple minimum-energy paths between the system’s two metastable states.
Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space. The method identifies multiple physically plausible pathways between the states, also observed with long MD simulations. For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions.

### v5

@@ Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states. These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling. We demonstrate our approach on two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms. Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study. Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in less than 15 minutes on a standard GPU rather than weeks on specialized hardware.
@@ 1 Introduction
Characterizing molecular conformational changes, or transitions, is essential for understanding biological processes, from enzyme catalysis to protein folding. Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes. This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935). As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive.
@@
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as splines or as a discrete series of molecular configurations connected by artificial spring forces to ensure continuity. Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for small-scale systems, it has not yet benefited from modern optimization approaches that could address these challenges.
@@
Recent advances in machine learning methodology have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical force field models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives beyond forces, which are needed for neural path optimization.
Numerous approaches have emerged over the past decades to tackle the problem of sampling transitions. Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics to begin with. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Lee et al., 2025; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
@@
Notably, Du et al. (2024) uses a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024). Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, a class of methods related to string methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics to optimize the neural path. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a force field to generate plausible transition paths using the discretized Onsager-Machlup action directly as a loss function. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to much larger biomolecular systems with explicit solvent, essential for accurately modeling such systems, while maintaining fast training times.
Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy-path finding problem as a continuous optimization framework. Our approach transforms the Onsager-Machlup functional, which naturally models the probability of paths in stochastic systems, into a remarkably simple and intuitive loss function that requires only energy evaluations along the neural network parametrized path. In contrast to existing methods that use complex composite losses involving energies, forces, and path gradients, our streamlined formulation eliminates the need for these additional computations. This simplification, while faster, also sidesteps known issues in PINN optimization if the loss contains differential operators (Rathore et al., 2024). This PINN framework efficiently captures physically meaningful transition mechanisms for complex biomolecular systems, completing optimization in minutes on standard hardware.
@@ 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To derive our loss function, which when minimized yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods. However, we further simplify the approach after reaching their geometric formulation. A complete derivation with all mathematical details is provided in Appendix A.1. We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
The exponent corresponds to the Onsager-Machlup action functional Onsager & Machlup (1953):
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. Notably, there are different derivations of the action that yield an additional higher-order derivative term related to trajectory entropy, whose necessity is conditional on the task (Adib, 2008). Since we are primarily interested in MEPs, we neglect this term. However, the resultant MEPs still empirically align well with the free energy landscape. To further simplify, we expand the integrand of the action functional:
@@
Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization. We can therefore focus on minimizing the remaining terms. Applying the Cauchy-Schwarz inequality to the remaining terms:
This upper bound becomes an equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape. Crucially, we allow the path to optimize its shape and length in our formulation. This freedom enables the path to adjust to follow the energy gradient (or its opposite), making the above upper bound tight at the optimum. When we abandon fixed-time parameterization and allow the path to stretch as needed to align with energy gradients, we can transform the time integral into a purely geometric one by defining the arc-length element along the spatial curve :
While the referenced works (Vanden-Eijnden & Heymann, 2008; Olender & Elber, 1997) derive and use this geometric action, we further simplify it. This geometric integral represents a natural endpoint for theoretical development for traditional chain-of-states and string methods. In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string using forces conceptually straightforward.
However, we want a loss function that acts directly on the neural network parameters in our neural network formulation. While we could parameterize the path using a neural network and then compute forces along the path before backpropagating to our parameters, this would introduce additional gradient computations, which have been noted to lead to optimization challenges in PINNs (Rathore et al., 2024). We avoid this issue by simplifying to an energy-based loss.
We represent the transition path as a continuous neural mapping from a progress coordinate to the system’s Cartesian coordinates, such that and . We then discretize the path into segments:
@@
For a path segment, applying Taylor’s theorem, we can express the energy difference as:
Where is the angle between the gradient and displacement vectors. For MEPs, the displacement aligns with or against the gradient direction () giving:
@@
Where is the lowest energy along the path. Since the latter terms are constants during optimization, this leads to our final loss function. To implement this in practice, we sample B points along a parametrized path and directly minimize the sum of energies at these points, yielding:
where are uniformly or randomly sampled points in . By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
2.1.1 Neural Representation of MEPs
We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms, as illustrated in Figure 1 (left):
Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the neural network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the neural network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
@@ 2.1.2 Path Optimization
We minimize using the Adam optimizer (Kingma & Ba, 2017):
Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary; the inherent continuity of , combined with smooth blending functions, ensures path continuity.
2.2 Established Approaches to MEP Finding
Classical approaches to MEP discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation. This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization. The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as by E & Vanden-Eijnden (2010). These established techniques illustrate both foundational concepts and limitations that our approach addresses.
@@ 2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized using splines mapping to the systems Cartesian coordinates that connects these states. The MEP satisfies:
@@
Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all . In practice, the string is discretized into images and evolved through a two-step procedure that requires explicit reparameterization steps:
@@ 2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition .
@@ 2.2.2 Chain-of-States Methods
An alternative perspective views MEP finding as optimization of a chain-of-states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008). Given a sequence with and , the objective function is defined as:
where and is a spring constant. The spring term enforces the continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost.
@@ 3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025). In Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, neural network architectures, and regularization approaches.
@@ 3.1 AIB9 Transition Path Discovery
We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods. This system effectively bridges abstract models and complex proteins: it is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes; its energy landscape is well-characterized; and its transitions are easily visualized through dihedral angle projections. Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields. We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space. The method identifies multiple physically plausible pathways between the states, which are also observed with long MD simulations. For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions. Notably, the discovered MEPs align well with the free energy landscape despite our choice of action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.